OpenStudy (anonymous):

If sin Θ = 1 over 3 and tan Θ < 0, what is the value of cos Θ

OpenStudy (dumbcow):

sin^2 + cos^2 = 1

OpenStudy (gorv):

\[\tan \theta =\frac{ \sin \theta }{ \cos \theta}\]

OpenStudy (gorv):

\[\sin \theta= \frac{ 1 }{ 3 }\]

OpenStudy (gorv):

use above 2 formula to find tan

OpenStudy (anonymous):



OpenStudy (anonymous):

@nbudel from the figure find the value of cos theta